arked out or 200P Hag question Determining Retalned Earnings and Net Income The following information appears in the records of Bock Corporation at year-end Accounts Receivable $23,000 Retained Earnings $7 Accounts Payable 11,000 Supplies Cash Common Stock 110,000 9,000 8,000 Equipment, net154,000 a. Calculate the balance in Retained Earnings at year-end. b. If the amount of the retained earnings at the beginning of the year was $30,000 and $12,000 in dividends is paid during the year, calculate net income for the year Save Answers Previous page Finish atten MacBook Air 4
Answer #1

Answers :

1. Ending Balance of Retained Earnings = $73,000

2. Net Income = $ 55,000

We Can solve this question With Accounting Equation

Accounting Equation :

Assets = Liabilities +Shareholders' Equity

Assets =Accounts Receivables + Cash+ Supplies + Equipment = (23,000+8,000+9,000+154,000)

Assets = $194,000

Liabilities = Accounts Payable = 11,000

Shareholders' Equity = Common Stock + Retained Earnings = 110,000 + ?

Retained Earnings = 194000-11,000-110,000 = $73,000

Retained Earnings Euqtion

BEginning RE + Net income - Dividned = Ending Retained Earnings

30,000+Net Income -12,000 = 73,000

Net income = $55,000